OpenStudy (adi3):

2m^2 * 4m^3/2 * 4m^-2

OpenStudy (adi3):

iget the that 2 * 4 * 4 = 32 but how to do the exponenets

OpenStudy (baru):

use this formula \(a^m\times a^n=a^{m+n}\)

OpenStudy (adi3):

i get -6

OpenStudy (baru):

for the exponent? how?

OpenStudy (baru):

\(32m^{2+\frac{3}{2}-2}\)

OpenStudy (adi3):